A chord 4.2cm long is 2.8cm from the centre of a circle. Calculate the radius of the circle.
Answer by Alan3354(69443)

A chord 4.2cm long is 2.8cm from the centre of a circle. Calculate the radius of the circle.
--------------
2.1^2 = (r + 2.8)*(r - 2.8)
4.41 = r^2 - 7.84
r^2 = 12.25
r = 3.5
===========